The term structure of interest rates is as follows:

  • 1-year rate = 5.5%
  • 2-year rate = 6.2
  • 3-year rate = 6.8

What does the market expect the one-year rate to be two years from today? Write your answer out to four decimals
Answer #1

Based on Pure Expectations Theory,

(1 + 3 Yr rate)3 = (1 + 2 Yr Rate)2 * (1 + 1 Yr Rate 2 Yr from now)

(1 + 6.8%)3 = (1 + 6.2%)2 * (1 + 1 Yr Rate 2 Yr from now)

1.218186 = 1.127844 * (1 + 1 Yr Rate 2 Yr from now)

(1 + 1 Yr Rate 2 Yr from now) = 1.080102

1 Yr rate 2 Yr from now = 0.080102 = 8.0102%